Use the distance formula to find the equation of a parabola with a focus at (0, -10) and a directrix as y = 10.
y = (1/40)x^2
y = x^2 + (1/40)
y = 40x^2
y = x^2 + 40

Answer by Boreal(15235)

Vertex is half way between focus and directrix.
That puts the vertex at 0.
p=10
x^2=4*10y
40y=x^2
y=(1/40)x^2